# Scaling properties of two-particle–two-hole responses in asymmetric nuclei for neutrino scattering within the Relativistic Mean-Field framework

### v1

@@ I Introduction
For simplicity, a universal coupling is assumed Weh93 , assigning the same energy shift to both nucleons and the $\Delta$ resonance, $E_{v}^{N}=E_{v}^{\Delta}$. The effective nucleon and $\Delta$ masses, together with the proton and neutron Fermi momenta employed for each nucleus, are summarized in Table 1. These parameters are chosen phenomenologically to reproduce inclusive electron-nucleus scattering data by extracting the 2p2h contribution from the quasielastic peak, varying the Fermi momentum and the effective nucleon mass to describe the experimental data, and subsequently adjusting the effective $\Delta$ mass to reproduce the $\Delta$-resonance region, using a superscaling-based methodology, further details of the fitting procedure can be found in Refs. Mar17 ; Mar21a .
@@ III Scaling of 2p2h Responses Relative to Carbon
In the $nn\to np$ channel (Fig. 3), we find the same overall trend, but with an absolute magnitude roughly six times smaller than in the $np\to pp$ channel. Here the scaling is visibly less accurate, consistent with the enhanced sensitivity of the $nn\to np$ response to neutron excess. The peak region is also broader and less sharply defined. This behavior follows from the interplay of $\Delta$ mechanisms: the $\Delta$-forward diagram dominates channels with two protons in the final state, whereas the $\Delta$-backward contribution becomes comparatively more relevant in the $np$ final channel, producing an effective smearing in the $\omega$ dependence, in line with the analysis in Ref. Mar24b . In both channels, $\mathcal{R}$ increases with atomic number, tracking the approximate proportionality of the 2p2h strength to the nuclear volume. In Appendix A, we compare the corresponding transverse electromagnetic $T$ responses for electron-induced reactions in the $nn$, npnp, and $pp$ channels, using coefficients different from those employed for neutrinos.
@@
In Figs. 4 and 5 we apply the same scaling analysis to the $T^{\prime}$ response in neutrino-induced reactions. The scaling factor is extracted from the $T$ response and then used for the remaining neutrino responses ($CC$, CLCL, LLLL, and $T^{\prime}$). In the energy range considered, $T$ and $T^{\prime}$ dominate the neutrino cross section. We also obtain the approximate factor $\sim 1/2$ between $T$ and $T^{\prime}$, as reported in Ref. Mar21b . Overall, the rescaled $T$ and $T^{\prime}$ responses follow the same pattern, with small channel-dependent differences. For example, in the $np$ final-state channel for 238U the maximum ratio extracted from $T^{\prime}$ is slightly larger than the reference value.
@@
A striking feature is the much steeper mass dependence of the $\nu(np)$ and $e(nn)$ channels. From helium to uranium, these ratios increase by a factor of about 325, whereas the remaining channels grow by $\sim 100$-$120$. This trend is quantified by the scaling ratios reported in Table 2, which span more than an order of magnitude across the nuclear chart. For example, in the neutrino-induced $np\to pp$ channel the ratio increases from $\mathcal{R}\simeq 0.185$ for 4He to $\mathcal{R}\simeq 18.5$ for 238U, whereas in the $nn\to np$ channel it reaches $\mathcal{R}\simeq 38.5$ for 238U (all values normalized to unity for 12C).
@@
48Ca 4.444 7.399 3.817 4.405 7.634
56Fe 4.049 5.625 4.016 4.149 5.376
59Ni 3.690 4.975 3.817 3.831 4.695
89Y 5.535 8.772 4.975 5.618 8.197
119Sn 7.657 12.987 6.369 7.634 12.195
181Ta 11.694 21.977 9.346 11.765 20.408
208Pb 14.01 27.316 10.753 13.889 25.641
238U 18.75 40.462 15.625 18.519 38.462
Figure 6: Transverse responses $R^{T}(X)$ rescaled to 12C by dividing by the corresponding scaling ratio $\mathcal{R}$, for the seventeen nuclei considered at $q=500$ MeV/$c$. Top: pppp, npnp, and $nn$ channels in electron scattering. Bottom: npnp (left) and pppp (right) final-state channels in neutrino scattering.
@@ IV Factorized Parametrization of the 2p2h Nuclear Response
$e+nn\rightarrow e+nn$ $0.51$ $0.41$ $1.82$ $0.73$
$e+np\rightarrow e+np$ 1.42-1.42 $0.35$ $1.89$ $0.94$
$e+pp\rightarrow e+pp$ $1.16$ 0.46-0.46 $1.74$ $0.97$
$\nu_{\mu}+nn\rightarrow\mu+np$ 0.78-0.78 $0.66$ $1.26$ $0.54$
$\nu_{\mu}+np\rightarrow\mu+pp$ 1.51-1.51 $0.56$ $2.11$ $0.98$
@@
Figure 8: Relative residuals as a function of the mass number AA for the five emission channels considered.
The values of $R^{2}$ (the coefficient of determination of the linear regression) reported in the table provide a global measure of the quality of the parametrization for each channel. As expected, the quality of the fit is not uniform across channels, reflecting the different degree of residual nuclear complexity once volume and leading phase-space effects are removed. In Fig. 7, we display the results of the global fit as a function of the mass number $A$. We observe that, for all channels, the factor $\rho$ tends to decrease with increasing nuclear size and approaches a nearly constant value for heavy nuclei. The figure also illustrates the relatively poor agreement obtained for the $nn$ channels in both electron and neutrino scattering, which contrasts with the very good description of the $pp$ channel. The latter is particularly relevant for neutrino-induced reactions, where it provides the dominant contribution.
@@
The resulting parametrization reproduces the main trends of the scaling factors across different nuclei and emission channels. As shown by the residual analysis, the relative residuals remain typically within $\pm 10\%$, corresponding to relative deviations of the order of $10\%$. As shown in Fig. 8, the relative residuals display the largest dispersion for the lightest nuclei, where sizable positive and negative deviations are observed. For intermediate-mass systems, the residuals cluster more tightly around zero, with most values lying within the $\pm 10$-15% band and exhibiting a smooth dependence on nuclear size. For the heaviest nuclei, the dispersion increases again, although it remains less pronounced than in the lightest systems and does not show a strong systematic bias.
@@ V Conclusions
The overall quality of the parametrization is moderate, capturing the main trends of the scaling ratios across the nuclear chart. For medium-mass nuclei, the relative deviations are typically at the $\sim 10\%$ level, as can be inferred from the residual analysis. Larger deviations are concentrated in the lightest and heaviest systems, reflecting both genuine nuclear-structure effects and the limitations imposed by the restricted set of microscopic calculations considered in this work, which spans 16 nuclei. These results illustrate that achieving a fully universal description across all nuclei and emission channels remains a challenging task within the present dataset.

### v2

@@ I Introduction
For simplicity, a universal coupling is assumed Weh93 , assigning the same vectorial energy to both nucleons and the $\Delta$ resonance, $E_{v}^{N}=E_{v}^{\Delta}$. The effective nucleon and $\Delta$ masses, together with the proton and neutron Fermi momenta employed for each nucleus, are summarized in Table 1. These parameters ($k_{Fp}$, $k_{Fn}$, $m_{N}^{*}$, and $m_{\Delta}^{*}$) are determined through a phenomenological fit to more than 20000 data points of inclusive electron-nucleus scattering available in the Quasielastic Electron Nucleus Scattering Archive of the University of Virginia archive ; archive2 . The resulting values have been further validated against recent high-precision measurements from JLab Mur19 and MAMI Mih24 . The parameters listed in Table 1 are consistent with those obtained in previous Superscaling analyses Mar17 ; Mar21a , which employed a single effective Fermi momentum. In this work, we use independent Fermi momenta for protons ($k_{Fp}$) and neutrons ($k_{Fn}$) to account for nuclear asymmetry which is equivalent to the value taken in the previous references. In our procedure, the Fermi momentum and the nucleon effective mass are varied to best describe the quasielastic peak in kinematics where the peak is clearly distinguishable (excluding regions of very low or high momentum transfer). Subsequently, the effective $\Delta$ mass $m_{\Delta}^{*}$ is adjusted to reproduce the $\Delta$-resonance maximum, ensuring the peak position remains consistent with Relativistic Fermi Gas predictions Mar21a .
@@ III Scaling of 2p2h Responses Relative to Carbon
In the $nn\to np$ channel (Fig. 3), we find the same overall trend, but with an absolute magnitude roughly six times smaller than in the $np\to pp$ channel. Here the scaling is visibly less accurate, consistent with the enhanced sensitivity of the $nn\to np$ response to neutron excess. The peak region is also broader and less sharply defined. This behavior follows from the interplay of $\Delta$ mechanisms: the $\Delta$-forward diagram dominates channels with two protons in the final state, whereas the $\Delta$-backward contribution becomes comparatively more relevant in the $np$ final channel, producing an effective smearing in the $\omega$ dependence, in line with the analysis in Ref. Mar24b . In both channels, $\mathcal{R}$ increases generally with atomic number, tracking the approximate proportionality of the 2p2h strength to the nuclear volume. In Appendix A, we compare the corresponding transverse electromagnetic $T$ responses for electron-induced reactions in the $nn$, npnp, and $pp$ channels, using coefficients different from those employed for neutrinos.
@@
In Figs. 4 and 5 we apply the same scaling analysis to the $T^{\prime}$ response in neutrino-induced reactions. The scaling factor is extracted from the $T$ response and then used for the remaining neutrino responses ($CC$, CLCL, LLLL, and $T^{\prime}$). In the energy range considered, $T$ and $T^{\prime}$ dominate the neutrino cross section. We also obtain the approximate factor $\sim 1/2$ between $T$ and $T^{\prime}$, as reported in Ref. Mar21b . Overall, the rescaled $T$ and $T^{\prime}$ responses follow the same pattern, with small channel-dependent differences. For example, in the $pp$ final-state channel for 208Pb the maximum ratio extracted from $T^{\prime}$ is slightly larger than the reference value.
@@
A striking feature is the much steeper mass dependence of the $\nu(np)$ and $e(nn)$ channels. From helium to uranium, these ratios increase by a factor of about 300, whereas the remaining channels grow by $\sim 100$-$120$. This trend is quantified by the scaling ratios reported in Table 2, which span more than an order of magnitude across the nuclear chart. For example, in the neutrino-induced $np\to pp$ channel the ratio increases from $\mathcal{R}\simeq 0.185$ for 4He to $\mathcal{R}\simeq 18.5$ for 238U, whereas in the $nn\to np$ channel it reaches $\mathcal{R}\simeq 38.5$ for 238U (all values normalized to unity for 12C).
@@
48Ca 4.210 6.881 3.484 4.172 6.976
56Fe 4.050 5.508 3.914 4.082 5.251
59Ni 3.755 5.011 3.759 3.798 4.636
89Y 5.389 8.457 4.772 5.482 7.874
119Sn 7.288 12.32 5.950 7.331 11.39
181Ta 11.10 20.03 8.442 11.12 18.65
208Pb 13.24 24.73 9.758 13.07 23.22
238U 17.90 36.79 14.48 17.76 35.95
Figure 6: Transverse responses $R^{T}(X)$ rescaled to 12C by dividing by the corresponding scaling ratio $\mathcal{R}$, for the seventeen nuclei considered at $q=500$ MeV/$c$. The dashed black lines represent the results for 12C, while the different colored lines correspond to the various rescaled nuclei considered in this work. Top: pppp, npnp, and $nn$ channels in electron scattering. Bottom: npnp (left) and pppp (right) final-state channels in neutrino scattering.
@@ IV Factorized Parametrization of the 2p2h Nuclear Response
$e+nn\rightarrow e+nn$ $0.77$ 0.09-0.09 $1.84$ $0.95$
$e+np\rightarrow e+np$ 0.58-0.58 0.20-0.20 $1.95$ $0.99$
$e+pp\rightarrow e+pp$ $0.90$ 0.20-0.20 $1.78$ $0.97$
$\nu_{\mu}+nn\rightarrow\mu+np$ $0.43$ 0.28-0.28 $1.34$ $0.96$
$\nu_{\mu}+np\rightarrow\mu+pp$ 0.64-0.64 0.13-0.13 $2.03$ $0.99$
@@
Figure 8: Relative residuals as a function of the mass number AA for the five emission channels considered. The purple circles represent the different nuclei analyzed in this work. The green shaded area indicates a relative residual interval of $\pm 10\%$, while the yellow shaded area corresponds to a $\pm 20\%$ relative residual interval.
The values of $R^{2}$ (the coefficient of determination of the linear regression) reported in the table provide a global measure of the quality of the parametrization for each channel. The results demonstrate a uniformly high quality of the fit accoss all channels, with $R^{2}$ above 0.95. In Fig. 7, we display the results of the separated channel fit as a function of the mass number $A$. We observe that, for all channels, the factor $\rho$ tends to decrease with increasing nuclear size and approaches a nearly constant value for heavy nuclei. The figure also illustrates the relatively very good description in all channel. The $pp$ channel with $R^{2}=0.99$, is particularly relevant for neutrino-induced reactions, where it provides the dominant contribution.
@@
The resulting parametrization reproduces the main trends of the scaling factors across different nuclei and emission channels. As shown by the residual analysis, the relative residuals remain typically within $\pm 10\%$, corresponding to relative deviations of the order of $10\%$. As shown in Fig. 8, the relative residuals display the largest dispersion for the lightest nuclei, where sizable positive and negative deviations are observed. For intermediate-mass systems, the residuals cluster more tightly around zero, with most values lying within the $5\%$ band and exhibiting a smooth dependence on nuclear size. For the heaviest nuclei, the dispersion increases again, although it remains less pronounced than in the lightest systems and does not show a strong systematic bias.
@@ V Conclusions
The overall quality of the parametrization is moderate, capturing the main trends of the scaling ratios across the nuclear chart. For medium-mass nuclei, the relative deviations are typically bellow the $\sim 10\%$ level, as can be inferred from the residual analysis. Larger deviations are concentrated in the lightest and heaviest systems, reflecting both genuine nuclear-structure effects and the limitations imposed by the restricted set of microscopic calculations considered in this work, which spans 16 nuclei. These results illustrate that achieving a fully universal description across all nuclei and emission channels remains a challenging task within the present dataset.